A coil of area 10-3 m2 and consisting of 30 turns of wire is placed with its plane at right angles to a magnetic field of flux density 0.1 T. When the coil is withdrawn from the field a meter in series with coil indicates that a charge of 100 mC has flown around the circuit. What is the combined resistance of the coil and meter?
1
Expert's answer
2020-09-14T12:53:36-0400

As per the given question,

Area of the coil "(A)=10^{-3}m^2"

Number of turns in the coil "(n)=30"

Flux density of the magnetic field =0.1 T

Charge "(q)=100mC"

Let the resistance in the coil is R=?

Change in flux "d\phi = N \Delta BA =30\times 0.1\times 10 =30 Tm^2"

We know that "q=\dfrac{d\phi}{R}"

Now, substituting the values,

"R=\frac{d\phi}{q}"

"\Rightarrow R=\dfrac{30 Tm^2}{100\times 10^{-3}C}"

"\Rightarrow R=300 \Omega"